If I am four times older than my daughter and in 20 years time I shall be twice as old as her. How old are we today?
Answer by richwmiller(17219)

f=4*d
f+20=2*(d+20)
f+20=2*d+40
f=2*d+20
4*d=2*d+20
2d=20
d=10
f=40
check
40+20=2(10+20)
60=60
ok